5. Profitability ratios Profitability ratios help in the analysis of the combined impact of liquidity ratios, asset management ratios, and debt management ratios on the operating performance of a firm Your boss has asked you to calculate the profitability ratios of Petroxy Oil Co. and make comments on its second-year performance as compared to its first-year performance The following shows Petroxy Oil Co.s income statement for the last two years. The company had assets of $4,700 million in the first year and $7,518 million in the second year. Common equity was equal to $2,500 million in the first year, and the company distributed 100% of its earnings out as dividends during the first and the second years. In addition, the firm did not issue new stock during either year Petroxy Oil Co Income Statement For the Year Ending on December 31 (Millions of dollars) Year 2 Yar 1 2,000 1,268 80 1,348 652 68 584 234 350 Net Sales 2,540 1,365 127 1,492 1,048 105 943 377 566 Operating costs except depreciation and amortization Depreciation and amortization Total Operating Costs Operating Income (or EBIT) Less: Interest Earnings before taxes (EBT) Less: Taxes (40%) Net IncomeCalculate the profitability ratios of Petroxy Oil Co. in the following table. Convert all calculations to a percentage rounded to two decimal places. Ratio Value Year 2 Year 1 Operating margin Profit margin Return on total assets Return on common equity Basic earning power 32.60% 22.28% 7.45% 14.00% 13.94% Decision makers and analysts look deeply into profitability ratios to identify trends in a companys profitability. Profitability ratios give insights into both the survivability of a company and the benefits that shareholders receive. Identify which of the following statements are true about profitability ratios. Check all that apply. □ If a company has a profit margin of 10%, it means that the company earned a net income of $0.10 for each dollar of sales. □ An increase in the return on assets ratio implies an increase in the assets a firm owns If a companys operating margin increases but its profit margin decreases, it could mean that the company paid more in interest or taxes. If a company issues new common shares but its net income does not increase, return on common equity will increase
Answer #1

Please provide rating...

answer 1 Year 2 Year 1
a Net sales 2540 2000
b Operating cost except depreciation and amortization 1365 1268
c Depreciation and amortization 127 80
d Total operating cost 1492 1348
e Operating income(EBIT) 1048 652
f Less : Interest 105 68
g EBT 943 584
h Less : Taxes 377 234
i Net income 566 350
j Total asset 7518 4700
k Equity 2500 2500
l=e/a Operating margin 41.26% 32.60%
m=i/a Profit margin 22.28% 17.50%
n=i/j Return on Total asset 7.53% 7.45%
o=i/k Return on total equity 22.64% 14.00%
p=e/j Basic earner power 13.94% 13.87%
answer 2 Correct statement is :
IF a company’s operating margin increases but its profit margin decreases , it could mean that the company paid more in interest or taxes